OpenStudy (anonymous):

In order for cos(x) = sin(x + h), h must be equal to: Is it 120, 60, 45, or 90?

OpenStudy (anonymous):

I know its one of those values

OpenStudy (ash2326):

\[\sin (\theta + 90)= \cos (\theta)\]

OpenStudy (anonymous):

f(x) is translated by vector\[\left| a \\ 0 \right|\] then it becomes f(x-a) since sin(x) is a translation of cos(x) by the vector where a= -90 we get the above answer

OpenStudy (anonymous):

or think about it like this we know that cos(x) = cos(-x) and we know that sin(90-x) = cos(x) from this triangle  so sin(90-x) = cos(x) = cos(-x) now let -x = y and we have cos(y) = sin(90 + y)

OpenStudy (ash2326):

\[\cos x=\sin (90-x)\] or \[ \cos x= \sin (90+x)\] here we have \[ \cos x= \sin (x+h)\] so here h= 90

OpenStudy (anonymous):

Thanks a lot guys:). I have to submit my assignment in 25 minutes. You guys saved my life!

OpenStudy (anonymous):

:D no problem happy to be of assistance